A rectangle with a length of x+4 and width of x+3, what is the perimeter of the rectangle if its area is 20cm^2?

Thanks

Answer by Boreal(15235)

area is x^2+7x+12=20
so x^2+7x-8=0
(x+8)(x-1)=0
x=1 cm only positive root.
the sides are 4 cm and 5 cm
The perimeter is 18 cm.